OpenStudy (anonymous):

3 over 4 x + 5 over 8 = 4x

OpenStudy (shamil98):

\[\large \frac{ 3 }{ 4 }x +\frac{ 5 }{ 8 } = 4x\] this?

OpenStudy (anonymous):

Yes!

OpenStudy (shamil98):

Okay. Get x on one side and the constant on the other. \[\large -3\frac{ 1 }{ 4 } x= -\frac{ 5 }{ 8 }\] Multiply by reciprocal to get rid of the fraction on x. \[\large x = -\frac{ 5 }{ 8 } \times -\frac{ 8 }{ 13 }\] \[\large x = \frac{ 5 }{ 13 }\]

OpenStudy (anonymous):

ok. but i'm kind of confused